At Zoom University, people's faces appear as circles on a rectangular screen. The radius of one's face is directly proportional to the square root of the area of the screen it is displayed on. Haydn's face has a radius of 2 on a computer screen with area 36. What is the radius of his face on a 16 × 9 computer screen?

Answer by Theo(13342)

the direct variation formula is y = kx
k is the constant of variation.
y is the radius of his face.
x is the square root of the area of the screen.
when y = 2 and the square root of 36 is 6, the formula becomes:
2 = k * 6
solve for k to get:
k = 2/6 = 1/3
when the screen size is 16 * 9, the area of the screen is 144.
the square root of the area of the screen is 12.
the formula becomes y = 1/3 * 12 = 4.
this says that the radius of his face is equal to 4 when the area of the screen is 144.
that sshould be your solution.